40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1853     Entry time: Mon Nov 2 18:23:46 2020
Author: Paco 
Type: DailyProgress 
Category: General 
Subject:  

Today; entered lab at ~09:08. I verified the orientations of the aspheric lenses and blaze gratings relative to the flextures, packaged and then dropped the parts for epoxying to Koji in 40m ~ 11:00. Spent some time between 12:00 and 12:45 finishing the ECDL connections. Everything looked good so I hooked it up to the TED200C controller. After a bit of research, I found out the Steinhart constants for the 10k thermistor;

a=1.129241e-3

b=2.341077e-4

c=8.775468e-8

Plugging these into the Steinhart equation give the actual temperatures from the Tact output on the TED200C (otherwise read as kOhm). According to the spec sheet, the TEC was tested at 250 mA (0.40 V), so not knowing a bunch more, set I_TEC on the TED200C to this limit and inspect the actual TEC current by scanning the Tset (setpoint) and recording the current in the ~ 15 - 25 deg C (attached plot, horizontal line marks room temperature). The diode current driver is hooked up, and everything is on the table as is. Left Crackle ~ 18:30.

Attachment 1: 2020_11_02_ted200c_test.png  50 kB  Uploaded Tue Nov 3 09:08:48 2020  | Hide | Hide all
2020_11_02_ted200c_test.png
ELOG V3.1.3-